OpenStudy (anonymous):

Integral (e^x + e^-x)^2 dx Solve the integral, What should I choose for u subsitution?

OpenStudy (zarkon):

multiply it out

OpenStudy (anonymous):

okay let me check it out.. thanks

OpenStudy (anonymous):

Yea im still stuck after that.. I'm getting (That 2nd term is supposed to be (e^(-x)) \[\int\limits_{}^{}(e^x)^2+2(e^x)(e^-x))+(e^-x))^2\]

OpenStudy (zarkon):

\[(e^x)^2=e^{2x}\]

OpenStudy (anonymous):

note that: \[e^{x}e^{-x} = e^{x-x} = e^0 = 1\]

OpenStudy (anonymous):

can you help me a little more, So I understand the algebra to make that equation but what u would i use to substitute

OpenStudy (anonymous):

1/2e^2x+2-1/2e^-2x+c

OpenStudy (anonymous):

dont sub. just evaluate

OpenStudy (zarkon):

you are missing a 2x

OpenStudy (anonymous):

I have: \[\int\limits1+2e^xe^-x dx\]

OpenStudy (zarkon):

e^{-x} \[e^{-x}\] ;)

OpenStudy (anonymous):

plus 2x you are correct z

OpenStudy (anonymous):

After you multiply it out, you want to break up the integral to see it easier: \[\int\limits_{}^{}e^{2x}+2+e^{-2x}dx = \int\limits e^{2x}dx+\int\limits2dx+\int\limits e^{-2x}dx\]

OpenStudy (anonymous):

Ahhh, I'm going to do a u subsitution for 2x and then make du = 2dx then evaluate \[\int\limits(e^u+e^{-u})du\]

OpenStudy (zarkon):

looks like you are missing some stuff there.

OpenStudy (anonymous):

hmm, well I have \[\int\limits(e^{2x}+2+e^{-2x})dx\] Then I set u = 2x , solve for du =2dx, and rewrite the integral, as i had above. Does that work?

OpenStudy (zarkon):

\[\int\limits(e^u+2+e^{-u})\frac{du}{2}\]

OpenStudy (anonymous):

Where did the 1/2 come from?

OpenStudy (zarkon):

\[du=2dx\] so \[dx=\frac{du}{2}\]

OpenStudy (anonymous):

Hmmm Im pretty lost right now.. Bnut got it correct up there (he was missing a 2x though ) Can you walk me through how to get there?

OpenStudy (zarkon):

\[\int\limits(e^{2x}+2+e^{-2x})dx\] let \[u=2x\] \[du=2dx\Rightarrow dx=\frac{du}{2}\] \[\int\limits(e^{u}+2+e^{-u})\frac{du}{2}\] \[=\frac{1}{2}\int\limits(e^{u}+2+e^{-u})du\] \[=\frac{1}{2}(e^u+2u-e^{-u})+c\] \[=\frac{1}{2}(e^{2x}+2(2x)-e^{-2x})+c\] \[=\frac{1}{2}e^{2x}+2x-\frac{1}{2}e^{-2x}+c\]

OpenStudy (anonymous):

Bravo! Thanks

OpenStudy (zarkon):

np..night ;)